OpenStudy (anonymous):

Hi everyone! If you are trying to find the domain of a function sin(xy)/(x^2+y^2-4), I know that x^2+y^2 can't = 4 but can someone tell me the proper way to express the domain? Thanks!

OpenStudy (abb0t):

Domain is all values for which the function is undefined. I think that your domain for the given function is all real, where (x,y) ≠ 0. But, if you use the (x,y) = (0,0) for your example, that makes the function: \(\sf \color{}{\frac{(sin(0))}{0}}\) = 0 and so domain is \(\mathbb{R}^2\) I believe

OpenStudy (abb0t):

I hope that makes sense. I am not good at explaining this type of domain stuff. I just kind of learned how to find it. Haha.

OpenStudy (anonymous):

well if x=2 or -2 while y=0 or vice versa, then that's a no go, so since I clearly can't have those conditions, do you still think my domain is really All Real Numbers?

OpenStudy (abb0t):

I didn't see the -4 at the end. Sorry. In that case, your domain is \(\sf \color{}{\mathbb{R}^2: x ≠ \pm ~2}\) I am not sure if notation is right.

OpenStudy (anonymous):

so since y isn't on the x-axis, when we speak about finding the domain, do we kinda just exclude it from mention?

OpenStudy (anonymous):

so maybe R2: X can't = + or - 2, y can't = 0 ???

OpenStudy (anonymous):

or would you just leave the y out of the conversation?

OpenStudy (amistre64):

the limit of a function does not depend on the value of the function at the point in question .... therefore a function can be discontinuous and still have a limit. But the domain is still restricted.

OpenStudy (abb0t):

@amistre64 always swaggin' it up :)

OpenStudy (amistre64):

x^2+y^2-4 \(\ne\) 0 x^2+y^2 \(\ne\) 4 x^2+2xy+y^2 \(\ne\) 4+2xy (x+y)^2 \(\ne\) 4+2xy x+y \(\ne\) \(\pm\sqrt{4+2xy}\) prolly a bad approach :) y^2 \(\ne\) 4-x^2 y \(\ne \pm\sqrt{4-x^2}\)

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

\[D=\{x,y:x\in R~and~y\ne\pm\sqrt{4-x^2}\}\] might be a good notation

OpenStudy (anonymous):

wow! thanks! :o)

OpenStudy (amistre64):

hope it makes sense :)

OpenStudy (anonymous):

kinda confused about why I have to mention "y" because the domain is along the x-axis after all...kinda confused about that

OpenStudy (amistre64):

to domain is in the xy plane

OpenStudy (amistre64):

a function of 2 variables; neither one is dependant on the other ... but there are certain combinations that provide bad math spots.

OpenStudy (anonymous):

oh yeah...i see that...i'm so used to playing with lines that I forgot I'm now playing with partials :o)

OpenStudy (amistre64):

could be: x,y in R : y ne ...

OpenStudy (anonymous):

could you re-post your final notation with how you would say it exatcly in words so I understand how to tell someone what that notation means?

OpenStudy (amistre64):

the domain is the set of all ordered pairs (x,y) in R^2, such that y does not equal (this) of (that).

OpenStudy (amistre64):

edit ... *(this) or (that)

OpenStudy (anonymous):

that's what that fancy notation means from up above with the big bold "D" ?

OpenStudy (amistre64):

essentially, yes. But i was using my "final" setup, not my initial idea

OpenStudy (anonymous):

D={x,y:x∈R and y≠±4−x2−−−−−√}

OpenStudy (anonymous):

thanks amistre!

OpenStudy (amistre64):

D = {......} ; the domain is the set of D = { (x,y) \(\in\) R^2 : .... } all ordered pairs (x,y) in R^2 such that ..... D = { (x,y) \(\in\) R^2 : y \(\ne\) _____ } y is not equal to what we found